OpenStudy (anonymous):

what is the axis of symmetry for the equation y=2x^2 +4x−8

OpenStudy (anonymous):

this equation isn't symmetric is there a typo?

OpenStudy (zpupster):

no it does have symmetry https://www.desmos.com/calculator/otro1dtaoo

OpenStudy (cwrw238):

this is a parabola and has symmetry

OpenStudy (zpupster):

For a quadratic function in standard form, y = ax2 + bx + c, the axis of symmetry is a vertical line -b/2a

OpenStudy (cwrw238):

y=2x^2 +4x−8 = 2 (x^2 + 2x - 4) now convert the parentheses to the appropriate form) or zpupster' s method is fine

OpenStudy (cwrw238):

shoot - i can't remember what the form (x + a)^2 - b is called!!!! zpupster?

OpenStudy (zpupster):

vertex form -idk